OpenStudy (anonymous):

Evaluate each expression under the given conditions. tan(2x); cos(x) = 3/5 x, in Quadrant 1

OpenStudy (anonymous):

use "double angle" formula for tangent.

OpenStudy (anonymous):

\[\tan(2\theta)=\frac{2\tan(\theta)}{1-\tan^2(\theta)}\] and you know that since \[\cos(\theta)=\frac{3}{5}\] then \[\tan(\theta)=\frac{4}{3}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

replace \[\tan(\theta)\] in the formula by \[\frac{4}{3}\] to get your answer